On a test, 5 students receive a 70%, 2 students receive a 90%, and 1 student receives an 85%. What is the average test score for the group of students?
Found 2 solutions by addingup, Boreal:
Answer by addingup(3677)

(5*70+2*90+85)/8 = 76.88

Answer by Boreal(15235)

5*0.70=3.50
2*0.90=1.80
1*0.85=0.85
8 students, 6.15 total
0.76875 or 77% ANSWER
5 are 6 7/8 below, or 34 3/8 below
1 is 8 1/8 above
2 are 13 1/8 above or 26 1/4